OpenStudy (anonymous):

Can someone please help me? -8x + 4y = 20

OpenStudy (bmmoffattb):

Yes. Hold on.

OpenStudy (bmmoffattb):

-8x+4y=20 can be simplified to make this easier.

OpenStudy (bmmoffattb):

We can divide the numbers all by 2 and get: -4x+2y=10

OpenStudy (bmmoffattb):

We can divide it by 2 again and get: -2x+y=5

OpenStudy (anonymous):

ok

OpenStudy (bmmoffattb):

Now we should put 1 number on one side. Preferably Y.

OpenStudy (bmmoffattb):

So we do this: -2x+y=5 +2x on this side and +2x on the other

OpenStudy (bmmoffattb):

As a result, this turns into: =y=2x+5

OpenStudy (anonymous):

Thank you so much for your help

OpenStudy (bmmoffattb):

No problem.

OpenStudy (anonymous):

I finally understand it now :)

OpenStudy (bmmoffattb):

Great! If you need any more help, I'm here.

OpenStudy (anonymous):

ok thank you, your help is much appreciated